OpenStudy (anonymous):

What is i times -i? I looked it up and it's 1, but how?

OpenStudy (zzr0ck3r):

\(i*-i=-i*i=-\sqrt{-1}\sqrt{-1}=-(-1)=1\)

OpenStudy (zzr0ck3r):

\(\sqrt{a}\sqrt{a}=a\)

OpenStudy (michele_laino):

by definition we have: \[i ^{2}=-1\] and please note that: \[i*(-i)=-i ^{2}\] so?